OpenStudy (anonymous):

The product of two consecutive positive integers is 1 more than their sum. Find the integers

OpenStudy (anonymous):

x(x+1)=x+(x+1)+1 solve

OpenStudy (anonymous):

3.561552813, -0.561552813.....i got this but it looks wrong

OpenStudy (experimentx):

let x and x+1 be your two integers the product is x(x+1) and the sum is x + (x+1) now apply condition as stated by question and solve.

OpenStudy (anonymous):

x(x + -1) = x + (x + 1) + 1 Reorder the terms: x(-1 + x) = x + (x + 1) + 1 (-1 * x + x * x) = x + (x + 1) + 1 (-1x + x2) = x + (x + 1) + 1 Reorder the terms: -1x + x2 = x + (1 + x) + 1 Remove parenthesis around (1 + x) -1x + x2 = x + 1 + x + 1 Reorder the terms: -1x + x2 = 1 + 1 + x + x Combine like terms: 1 + 1 = 2 -1x + x2 = 2 + x + x Combine like terms: x + x = 2x -1x + x2 = 2 + 2x

OpenStudy (anonymous):

is that being done right so far

OpenStudy (anonymous):

thankx